A car travelling at 60 km /hr overtakes another car travelling at 42 km/hr assuming each car to be 5.0m long ,find the time taken during the overtake and thetotal road distance used for overtake

Grade:11

1 Answers

inam ullah
6 years ago
V(a)= 60km/hV(b) = 42km/hA and b are same direction so relatives velocity is = 60- 42=18Distances=S=5m S=Vtt= S/V t= 5/18t= .27